find the dimensions of a rectangle if its diagonal is 17 meters and its perimeter is 46 meters
Found 2 solutions by Alan3354, addingup:
Answer by Alan3354(69443)

find the dimensions of a rectangle if its diagonal is 17 meters and its perimeter is 46 meters
------------
P = 2W + 2L = 46
W + L = 23
----
W^2 + L^2 = 17^2
Can you do the rest?

Answer by addingup(3677)

If I split the rectangle in 2 I have two right triangles, and the dividing line, which measures 17, is the hypotenuse for both triangles. The perimeter of each triangle will be 46/2= 23. Since the hypotenuse is 17, that leaves 23-17= 6 meters for two sides of the triangle. Impossible. Check your numbers, either your hypotenuse is 7 and not 17 or your perimeter is 146 and not 46 or something like that.